OpenStudy (a1234):

Use the Law of Cosines and find side b. Is it 15.38?

OpenStudy (a1234):

 a = 29, c = 39, and theta is 20 degs

OpenStudy (anonymous):

is this a right triangle?

OpenStudy (a1234):

No

OpenStudy (anonymous):

nvm was thinking something else...let me check

OpenStudy (campbell_st):

just substitute and calculate the value...

OpenStudy (anonymous):

you mean c^2=

OpenStudy (a1234):

I used the formula...my answer is in the question

OpenStudy (campbell_st):

oops \[b^2 = a^2 + c^2 - 2ac \cos(\theta)\]

OpenStudy (campbell_st):

so then you have \[b^2 = 29^2 + 39^2 - 2 \times 29 \times 39 \times \cos(20)\]

OpenStudy (campbell_st):

or you can just google a triangle solver, put in what you know and check your answer that way

OpenStudy (anonymous):

I also got 15.38